OpenStudy (anonymous):

Expand the logarithmic expression. log5^7a^5

OpenStudy (michele_laino):

hint: I can rewrite your expression as below: \[\log \left( {{5^7}{a^5}} \right) = \log \left\{ {\left( {{5^7}} \right) \cdot \left( {{a^5}} \right)} \right\}\] now, please apply this properties: \[\begin{gathered} \log \left( {{x^k}} \right) = k\log x,\quad k \in \mathbb{R} \hfill \\ \log \left( {x \cdot y} \right) = \log x + \log y,\quad x,y > 0 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

oops.. apply these* properties...